In each cycle of a Carnot engine, 97 J of heat is
absorbed from the high-temperature reservoir
and 50 J is exhausted to the low-temperature
reservoir.
What is the efficiency of the engine?
1
Expert's answer
2011-06-09T15:07:49-0400
This maximum efficiency η is defined to be:
η= W/QH = (T1-T2)/T1 =1-T2/T1 =1-Q2/Q1

Where W is the work done by the system (energy exiting the system as work), QH is the heat put into the system (heat energy entering the system), T2 is the absolute temperature of the cold reservoir, and T1 is the absolute temperature of the hot reservoir.
So
η=1-Q2/Q1 =1-50/97=1-0.515=0.485.